# Locally-APN Binomials with Low Boomerang Uniformity in Odd Characteristic

### v1

@@ 3.1 Differential Uniformity of $F_{r}$
Let $r>1$ be an integer and $q=p^{n}\equiv 3\pmod{4}$ If $(x+1)^{r}-x^{r}=b$ has at most $1$ solution in $S_{00}$ for all $b\in\mathbb{F}_{q}^{*}$ and $\gcd(r,q-1)\mid 2$, then $\delta_{F_{r}}(1,b)\leq 2$ for all $b\in\mathbb{F}_{q}^{*}$ and
@@
Note that if $x^{r}$ is PN, then the assumption of the above theorem holds clearly. The case $r=3$ will be discussed separately in Section 3.2, where we investigate the differential spectrum of $F_{3}$. We also find the differential spectrum of the case $r=\frac{2q-1}{3}$ in Section 3.3. To prove this theorem for the remaining cases, we investigate each $D_{ij}(b)$.
@@
$D_{11}(b)=\begin{cases}\frac{p^{n}-3}{4},&\text{ if }b=0,\\ 0,&\text{ otherwise.}\end{cases}$
@@
If $x\in S_{11}$, then (5) implies $b=0$. By Lemma 2.3, we obtain $S_{11}=\frac{p^{n}-3}{4}$, as desired. ∎
@@
By assumption, we can easily see that $D_{00}(b)\leq 1$. If $x^{r}$ is PN, then there is at most one such $x$ satisfying (7), we have $D_{00}(b)\leq 1$.
If there is a solution $x\in S_{00}$ in (5) with $b=0$, then we have $(x+1)^{r}=x^{r}$. If $\gcd(r,q-1)=1$, then we have $x=x+1$, a contradiction. If $\gcd(r,q-1)=2$, then we obtain $x^{2}=(x+1)^{2}$, which implies $x=-\frac{1}{2}$. Then, we get $\chi(x)=\chi\left(-\frac{1}{2}\right)=\chi\left(\frac{1}{2}\right)=\chi(x+1)=1$, a contradiction. Therefore, $D_{00}(0)=0$. ∎
@@
If there is a solution $x\in S_{01}$ of (5), then $x^{r}=-\frac{b}{2}$. Since $x\in S_{01}$, we have $\chi\left(\frac{b}{2}\right)=-1$. By the way, if there is a solution $x^{\prime}\in S_{10}$ of (5), then $(x^{\prime}+1)^{r}=\frac{b}{2}$. Since $x^{\prime}\in S_{10}$, we obtain $\chi\left(\frac{b}{2}\right)=-1$. Since $\chi\!\left(\frac{b}{2}\right)=1$ and $\chi\!\left(\frac{b}{2}\right)=-1$ cannot occur simultaneously, we conclude that $D_{01}(b)=0$ or $D_{10}(b)=0$. ∎
@@
By symmetry, the other proofs are similar with above, and we omit here. ∎
We have now analyzed all the cases of $D_{ij}(b)$, and are ready to prove the main theorem of this subsection for all $r$ except $r=3$ and $r=\frac{2q-1}{3}$. It is worth noting that the lemmas established here also hold for these two exponents, and hence we have (8), which will be proved in Lemma 3.6. These lemmas will also be used in the following subsections, where we investigate the differential spectra of $F_{r}$ for $r=3$ and $r=\frac{2q-1}{3}$.
@@
It is clear that $(x+1)^{r}-x^{r}=b$ has at most $1$ solution in $S_{00}$ for all $b\in\mathbb{F}_{q}^{*}$, when $x^{r}$ is PN. It remains to show that $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$ when $x^{r}$ is APN, that is, for $r\in\left\{\frac{3^{\frac{n+1}{2}}-1}{2},\ \frac{3^{n+1}-1}{8}\right\}$, where $p=3$ and $n$ is odd. In [7], where the APN property of $x^{r}$ was proved, the key idea was to show that the equation $(x+1)^{r}-x^{r}=b$ admits at most one solution in each of the two sets $S_{00}\cup S_{11}$ and $S_{01}\cup S_{10}$. Consequently, the condition $D_{00}(b)\leq 1$ follows directly from their argument. ∎
@@ 3.2 Differential Spectrum when $r=3$
We already observe that $\delta_{F_{3}}(1,0)=\frac{q+1}{4}$, in Section 3.1. By Lemma 3.2 and 3.4, we have
@@
If $b=\frac{1}{2}$, then (11) has one solution $x=-\frac{1}{2}$. But, we have $\chi(x+1)=\chi\left(\frac{1}{2}\right)\neq\chi\left(-\frac{1}{2}\right)=\chi(x)$ in this case, and hence $\frac{1}{2}\not\in S_{00}$. So, $D_{00}\left(-\frac{1}{2}\right)=0$. ∎
@@
implies that $\chi(2b)=-1$. Applying $\frac{2q-1}{3}$-th power on the both sides of (12), we have
@@
Since $\frac{2q-1}{3}$ is odd, we have $\chi(x)=\chi\left(-\frac{b}{2}\right)=-\chi(2b)=1$. Moreover,
@@
Next, we count the number of $b\in\mathbb{F}_{q}^{*}$ satisfying $\delta_{F_{3}}(1,b)=2$. By Lemma 3.6, we have $D_{00}(2)=0$, and hence $\delta_{F_{3}}(1,2)=1$, using (10). Hence, $\delta_{F_{3}}(1,b)=2$ if and only if $D_{00}(b)=1$ and $D_{01}(b)+D_{10}(b)=1$. So, by Lemmas 3.6 and 3.7, $\delta_{F_{3}}(1,b)=2$ if and only if $\chi(2b-1)=\chi(2b-4)=1$ and $\chi\left(b\left(b^{\frac{2q-1}{3}}-\alpha\right)\right)=-1$. Hence, we have
$8\omega_{2}=$ $-2+\displaystyle\sum_{x\in\mathbb{F}_{q}}(1+\chi(2x-1))(1+\chi(2x-4))(1-\chi\left(x\left(x^{\frac{q+1}{6}}-\alpha\right)\right))$
@@
where $S=\left\{x\in\mathbb{F}_{q}:\chi(2x^{3}-1)=\chi(2x^{3}-4)=1\right\}$. Thus,
@@
Since
@@
$=-1+\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi(u)\chi\left(u-\alpha\right)\chi(2u^{3}-1)+\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi(u)\chi\left(u-\alpha\right)\chi(2u^{3}-4)$
$+\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi(u)\chi\left(u-\alpha\right)\chi(2u^{3}-1)\chi(2u^{3}-4),$
$\omega_{2}\geq\frac{1}{8}\left(q-2-15\sqrt{q}\right)$, by Lemma 2.6. Hence, we can see that $\omega_{2}>0$, and hence $F_{3}$ is locally-APN if $q>231$, via SageMath. We confirm that $F_{3}$ is also locally-APN for $q<231$ via SageMath (see Table 4). ∎
@@ 3.3 Differential Spectrum when $r=\frac{2q-1}{3}$
Let $r=\frac{2q-1}{3}$ and $q\equiv 3\pmod{4}$. Then,
@@
is a square. If then, we obtain
@@
$x=X^{3}=\frac{\left(-3b^{2}\pm\left(3b(32-b^{3})\right)^{\frac{q+1}{4}}\right)^{3}}{1728b^{3}}=\frac{-36b^{2}\pm(4+b^{3})\left(3b(32-b^{3})\right)^{\frac{q+1}{4}}}{72b^{2}}.$
Hence, we can see that if we denote $x=x_{1}$ is a solution of (15), then the other is $x=x_{2}=-x_{1}-1$. If both $x_{1}$ and $x_{2}$ are in $S_{00}$, then $\chi(x_{1}+1)=\chi(-x_{2})=-1$ and $\chi(x_{2}+1)=\chi(-x_{1})=-1$, a contradiction. Hence, (15) has at most one solution in $S_{00}$. Also, one can check that
@@
$\omega_{2}$ $=\frac{1}{4}\left(\displaystyle\sum_{x\in\mathbb{F}_{q}}\left(1+\chi(x(32-x^{3}))\right)\left(1-\chi(x(x^{3}-8))\right)-3\right)$
$=\frac{1}{4}\left(\displaystyle\sum_{x\in\mathbb{F}_{q}}\left(1+\chi(x(32-x))\right)\left(1-\chi(x(x-8))\right)-3\right)$
$=\frac{1}{4}\left(q-3+\displaystyle\sum_{x\in\mathbb{F}_{q}}\chi(x(32-x))-\displaystyle\sum_{x\in\mathbb{F}_{q}}\chi(x(x-8))-\displaystyle\sum_{x\in\mathbb{F}_{q}}\chi(x^{2}(x-8)(32-x))\right)$
$=\frac{q-3}{4}$
Applying $\omega_{\frac{q+1}{4}}=1$ and $\omega_{2}=\frac{q-3}{4}$ on (14), we have $\omega_{1}=\frac{q+5}{4}$ and $\omega_{0}=\frac{q-3}{2}$. We complete the proof. ∎
@@ 4 Boomerang Properties
In this subsection, we study boomerang properties of $F_{r}$. We consider to find the number of common solutions $(x,y)$ of the following system.
@@ 4.1 Boomerang Uniformity of $F_{r}$
Let $b\in\mathbb{F}_{q}^{*}$. If $\gcd(r,q-1)\in\{1,2\}$, then there is no solution $(x,y)$ of (17) satisfying $x\in(S_{11}\cup\{-1\})$ or $y\in(S_{11}\cup\{-1\})$.
@@
By symmetry, we only show that there is no solution $(x,y)$ satisfying $x\in(S_{11}\cup\{-1\})$. Suppose that there is such a solution $(x,y)$ satisfying $x\in(S_{11}\cup\{-1\})$. If $y\in S_{00}$, then (17) reduces to $y^{r}=(y+1)^{r}=-\frac{b}{2}$. This impies that $y^{\gcd(r,q-1)}=(y+1)^{\gcd(r,q-1)}$. If $\gcd(r,q-1)=1$, then we have $y=y+1$, a contradiction. If $\gcd(r,q-1)=2$, then we obtain $y^{2}=(y+1)^{2}$, which implies $y=-\frac{1}{2}$. Then, we get $\chi(y)=\chi\left(-\frac{1}{2}\right)=\chi\left(\frac{1}{2}\right)=\chi(y+1)=1$, a contradiction.
@@
Let $b\in\mathbb{F}_{q}^{*}$. If
$\begin{cases}x^{r}-y^{r}=\frac{b}{2},\\ (x+1)^{r}-(y+1)^{r}=\frac{b}{2}.\end{cases}$ (18)
does not have a solution $(x,y)\in S_{00}\times S_{00}$, then $B_{0000}(b)=0$. Furthermore, if $D_{00}(b)\leq 1$ for all $b\in\mathbb{F}_{q}^{*}$, then $B_{0000}(b)=0$.
@@
If $(x,y)\in S_{00}\times S_{00}$, then (17) leads to (18). Then we have $x^{r}-y^{r}=\frac{b}{2}=(x+1)^{r}-(y+1)^{r}$, which implies $(x+1)^{r}-x^{r}=(y+1)^{r}-y^{r}$. By Lemma 3.3, $D_{00}(0)=0$ and hence $(x+1)^{r}-x^{r}=(y+1)^{r}-y^{r}=c$ for some $c\in\mathbb{F}_{q}^{*}$. Since $D_{00}(c)\leq 1$ for all $c\in\mathbb{F}_{q}^{*}$, we have $x=y$, and hence $b=0$, a contradiction. ∎
@@
The second equation of (19) has $\gcd(r,q-1)$ solutions. If $\gcd(r,q-1)=1$, then there exists a unique $x\in\mathbb{F}_{q}$ satisfying it. If $\gcd(r,q-1)=2$, then $r$ is even and hence $x^{r}=(-x)^{r}$ for all $x\in\mathbb{F}_{q}$. Thus, if $x=\theta$ is a solution of $(x+1)^{p^{k}+1}=\frac{b}{2}$, the other solution is $x=-\theta-2$. If both $\theta$ and $\theta-2$ lie in $S_{00}$, then $1=\chi(\theta+1)=\chi(-\theta-1)=-\chi(\theta+1)$, a contradiction. Therefore, the second equation of (19) admits at most one solution in $S_{00}$ in either case. For such $x$, the first equation of (19) has at most one solution $y$ with $\chi(y)=1$ in the same manner. Hence, $B_{0001}(b)\leq 1$, which completes the proof. ∎
@@
If (20) has a solution $y\in S_{01}$, then we have $2y^{r}=2=b$, which applies $y^{r}=-1$ and hence $r$ is odd. Then we obtain $\gcd(r,q-1)=1$ and y=1y=-1, which contradicts that $y\in S_{01}$.
@@
Combining Lemma 4.5 and Lemma 4.6 with the above equation, we have $\beta_{F}(1,b)\leq 2$, which completes the proof. ∎

### v2

@@ 3.1 Differential Uniformity of $F_{r}$
Let $r>1$ be an integer and $q=p^{n}\equiv 3\pmod{4}$. If $(x+1)^{r}-x^{r}=b$ has at most $1$ solution in $S_{00}$ for all $b\in\mathbb{F}_{q}^{*}$ and $\gcd(r,q-1)\mid 2$, then $\delta_{F_{r}}(1,b)\leq 2$ for all $b\in\mathbb{F}_{q}^{*}$ and
@@
Note that if $x^{r}$ is PN, then the assumption of the above theorem clearly holds. In this subsection, we prove Theorem 3.1 and verify its assumption for the exponents $r$ listed in Table 3. For the cases $r=3$ and $r=\frac{2q-1}{3}$, the verification of the assumption will be carried out separately in Sections 3.2 and 3.3, respectively, where we also determine the corresponding differential spectra. To handle the remaining cases, we analyze each $D_{ij}(b)$.
@@
$D_{11}(b)=\begin{cases}\frac{q-3}{4},&\text{ if }b=0,\\ 0,&\text{ otherwise.}\end{cases}$
@@
If $x\in S_{11}$, then (5) implies $b=0$. By Lemma 2.3, we obtain $S_{11}=\frac{q-3}{4}$, as desired. ∎
@@
By assumption, we can easily see that $D_{00}(b)\leq 1$.
If there is a solution $x\in S_{00}$ in (5) with $b=0$, then we have $(x+1)^{r}=x^{r}$. If $\gcd(r,q-1)=1$, then we have $x=x+1$, a contradiction. If $\gcd(r,q-1)=2$, then we obtain $x^{2}=(x+1)^{2}$, which implies $x=-\frac{1}{2}$. Then, we get $1=\chi(x)\chi(x+1)=\chi\left(-\frac{1}{4}\right)=\chi(-1)$, which is a contradiction since $\chi(-1)=-1$ when $q\equiv 3\pmod{4}$. Therefore, $D_{00}(0)=0$. ∎
If $x^{r}$ is PN, then there is at most one such $x$ satisfying (7), we have $D_{00}(b)\leq 1$.
@@
If there is a solution $x\in S_{01}$ of (5), then $x^{r}=-\frac{b}{2}$. Since $x\in S_{01}$, we have $\chi\left(\frac{b}{2}\right)=\chi(-x^{r})=-1$. By the way, if there is a solution $x^{\prime}\in S_{10}$ of (5), then $(x^{\prime}+1)^{r}=\frac{b}{2}$. Since $x^{\prime}\in S_{10}$, we obtain $\chi\left(\frac{b}{2}\right)=\chi\left((x^{\prime}+1)^{r}\right)=1$. Since $\chi\!\left(\frac{b}{2}\right)=1$ and $\chi\!\left(\frac{b}{2}\right)=-1$ cannot occur simultaneously, we conclude that $D_{01}(b)=0$ or $D_{10}(b)=0$. ∎
@@
By symmetry, the other proofs are similar to above, and we omit here. ∎
We have now analyzed all the cases of $D_{ij}(b)$, and are ready to prove the main theorem of this subsection for all $r$ except $r=3$ and $r=\frac{2q-1}{3}$. It is worth noting that the lemmas established here also hold for these two exponents, and hence we have (8), which will be proved in Lemmas 3.6 and 3.9. These lemmas will also be used in the following subsections, where we investigate the differential spectra of $F_{r}$ for $r=3$ and $r=\frac{2q-1}{3}$.
@@
Thus, the first part of Theorem 3.1 is proved, namely that $F_{r}$ is locally-APN with $\delta_{F_{r}}=\delta_{F_{r}}(1,0)=\frac{q+1}{4}$ under the given assumptions. We now turn to the second part, where we verify that the condition $(x+1)^{r}-x^{r}=b$ has at most one solution in $S_{00}$ for the exponents $r$ listed in Table 3. It is clear that $(x+1)^{r}-x^{r}=b$ has at most $1$ solution in $S_{00}$ for all $b\in\mathbb{F}_{q}^{*}$, when $x^{r}$ is PN. For $r\in\left\{\frac{3^{\frac{n+1}{2}}-1}{2},\ \frac{3^{n+1}-1}{8}\right\}$, it was shown in [7] that the equation $(x+1)^{r}-x^{r}=b$ admits at most one solution in each of the sets $S_{00}\cup S_{11}$ and $S_{01}\cup S_{10}$. Consequently, the condition $D_{00}(b)\leq 1$ follows directly from their argument. ∎
@@ 3.2 Differential Spectrum when $r=3$
We already observe that $\delta_{F_{3}}(1,0)=\frac{q+1}{4}$, in Section 3.1. By Lemmas 3.2, 3.4 and 3.5, we have
@@
If $b=\frac{1}{2}$, then (11) has one solution $x=-\frac{1}{2}$. But, we have $\chi(x+1)=\chi\left(\frac{1}{2}\right)\neq\chi\left(-\frac{1}{2}\right)=\chi(x)$ in this case, and hence $\frac{1}{2}\not\in S_{00}$. So, $D_{00}\left(\frac{1}{2}\right)=0$. ∎
@@
implies that $\chi(2b)=-1$. Since $\frac{2q-1}{3}$ is odd, raising both sides of (12) to the $\frac{2q-1}{3}$-th power, we have
@@
Consequently, we have $\chi(x)=\chi\left(-\frac{b}{2}\right)=-\chi(2b)=1$. Moreover,
@@
Next, we count the number of $b\in\mathbb{F}_{q}^{*}$ satisfying $\delta_{F_{3}}(1,b)=2$. By Lemma 3.6, we have $D_{00}(2)=0$, and hence $\delta_{F_{3}}(1,2)=1$, using (10). Hence, $\delta_{F_{3}}(1,b)=2$ if and only if $D_{00}(b)=1$ and $D_{01}(b)+D_{10}(b)=1$. Since $q\equiv 11\pmod{12}$, we have $\chi(3)=1$. Thus, by Lemmas 3.6 and 3.7, $\delta_{F_{3}}(1,b)=2$ if and only if $\chi(2b-1)=\chi(2b-4)=1$ and $\chi\left(b\left(b^{\frac{2q-1}{3}}-\alpha\right)\right)=-1$. Hence, we have
$8\omega_{2}=$ $\sum_{\begin{subarray}{c}x\in\mathbb{F}_{q}\\ x\neq 0,\,\frac{1}{2},\,2\end{subarray}}(1+\chi(2x-1))(1+\chi(2x-4))\left(1-\chi\left(x\left(x^{\frac{2q-1}{3}}-\alpha\right)\right)\right)$
$=$ $-2+\displaystyle\sum_{x\in\mathbb{F}_{q}}(1+\chi(2x-1))(1+\chi(2x-4))\left(1-\chi\left(x\left(x^{\frac{2q-1}{3}}-\alpha\right)\right)\right)$
@@
where $S=\left\{x\in\mathbb{F}_{q}:\chi(2x^{3}-1)=\chi(2x^{3}-4)=1\right\}$. Note that if $u^{3}=\frac{1}{2}$, then $1+\chi(2u^{3}-4)=1-\chi(3)=0$, and if $u^{3}=2$, then $u=\alpha$ and hence $\chi(u-\alpha)=0$, so these cases do not contribute to the sum. Thus,
@@
To complete the proof, it remains to show that $\omega_{2}>0$. Since
@@
$=-1+\Sigma_{1}+\Sigma_{2}+\Sigma_{3},$
where
$\Sigma_{1}$ $=\sum_{u\in\mathbb{F}_{q}}\chi\bigl(u(u-\alpha)(2u^{3}-1)\bigr),$
$\Sigma_{2}$ $=\sum_{u\in\mathbb{F}_{q}}\chi\bigl(u(u-\alpha)(2u^{3}-4)\bigr),$
$\Sigma_{3}$ $=\sum_{u\in\mathbb{F}_{q}}\chi\bigl(u(u-\alpha)(2u^{3}-1)(2u^{3}-4)\bigr).$
Since the corresponding polynomials have degrees $5,5,$ and $8$, respectively, and are not squares, Lemma 2.6 gives
$|\Sigma_{1}|\leq 4\sqrt{q},\qquad|\Sigma_{2}|\leq 4\sqrt{q},\qquad|\Sigma_{3}|\leq 7\sqrt{q}.$
Hence,
$D_{2}\leq-1+15\sqrt{q},$
and therefore
$\omega_{2}=\frac{1}{8}(q-3-D_{2})\geq\frac{1}{8}(q-2-15\sqrt{q}).$
In particular, $\omega_{2}>0$ for $q>231$.
For the remaining values $q<231$, we verify by direct computation (using SageMath) that $\omega_{2}>0$ (see Table 4). This completes the proof. ∎
@@ 3.3 Differential Spectrum when $r=\frac{2q-1}{3}$
Let $r=\frac{2q-1}{3}$ and $q\equiv 11\pmod{12}$. Then,
@@
is a square. Then, we obtain
@@
$x$ $=X^{3}=\frac{\left(-3b^{2}\pm\left(3b(32-b^{3})\right)^{\frac{q+1}{4}}\right)^{3}}{1728b^{3}}$
$=\frac{-27b^{6}\pm 27b^{4}\left(3b(32-b^{3})\right)^{\frac{q+1}{4}}-27b^{3}(32-b^{3})\pm 3b(32-b^{3})\left(3b(32-b^{3})\right)^{\frac{q+1}{4}}}{1728b^{3}}$
$=\frac{-864b^{3}\pm 24b(4+b^{3})\left(3b(32-b^{3})\right)^{\frac{q+1}{4}}}{1728b^{3}}=\frac{-36b^{2}\pm(4+b^{3})\left(3b(32-b^{3})\right)^{\frac{q+1}{4}}}{72b^{2}}.$
Hence, we can see that if we denote $x=x_{1}$ is a solution of (15), then the other is $x=x_{2}=-x_{1}-1$, since the two solutions satisfy $x_{1}+x_{2}=-1$. If both $x_{1}$ and $x_{2}$ are in $S_{00}$, then $\chi(x_{1}+1)=\chi(-x_{2})=-1$ and $\chi(x_{2}+1)=\chi(-x_{1})=-1$, a contradiction. Hence, (15) has at most one solution in $S_{00}$. Also, one can check that
@@
$4\omega_{2}$ $=\displaystyle\sum_{\begin{subarray}{c}x\in\mathbb{F}_{q}^{*}\\ x\neq 2,\,2\alpha^{2}\end{subarray}}\left(1+\chi(x(32-x^{3}))\right)\left(1-\chi(x(x^{3}-8))\right)$
$=\displaystyle\sum_{x\in\mathbb{F}_{q}^{*}}\left(1+\chi(x(32-x^{3}))\right)\left(1-\chi(x(x^{3}-8))\right)-2.$
Applying $u=x^{3}$ and since $\frac{2q-1}{3}$ is odd,
$4\omega_{2}$ $=\displaystyle\sum_{u\in\mathbb{F}_{q}^{*}}\left(1+\chi(u^{\frac{2q-1}{3}}(32-u))\right)\left(1-\chi(u^{\frac{2q-1}{3}}(u-8))\right)-2$
$=\displaystyle\sum_{u\in\mathbb{F}_{q}}\left(1+\chi(u(32-u))\right)\left(1-\chi(u(u-8))\right)-3.$
By Lemma 2.4,
$4\omega_{2}$ $=\displaystyle\sum_{u\in\mathbb{F}_{q}}1+\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi(u(32-u))-\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi(u(u-8))-\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi(u^{2}(u-8)(32-u))-3$
$=q-1-\displaystyle\sum_{u\in\mathbb{F}_{q}^{*}}\chi(u^{2}(u-8)(32-u))=q-1-\displaystyle\sum_{u\in\mathbb{F}_{q}^{*}}\chi((u-8)(32-u))$
$=q-2-\displaystyle\sum_{u\in\mathbb{F}_{q}}\chi((u-8)(32-u))=q-3.$
Hence, we obtain $\omega_{2}=\frac{q-3}{4}$. Applying $\omega_{\frac{q+1}{4}}=1$ and $\omega_{2}=\frac{q-3}{4}$ on (14), we have $\omega_{1}=\frac{q+5}{4}$ and $\omega_{0}=\frac{q-3}{2}$. We complete the proof. ∎
@@ 4 Boomerang Properties
In this subsection, we study boomerang properties of $F_{r}$. We consider the problem of finding the number of common solutions $(x,y)$ of the following system.
@@ 4.1 Boomerang Uniformity of $F_{r}$
Let $q\equiv 3\pmod{4}$ and $b\in\mathbb{F}_{q}^{*}$. If $\gcd(r,q-1)\in\{1,2\}$, then there is no solution $(x,y)$ of (17) satisfying $x\in(S_{11}\cup\{-1\})$ or $y\in(S_{11}\cup\{-1\})$.
@@
By symmetry, we only show that there is no solution $(x,y)$ satisfying $x\in(S_{11}\cup\{-1\})$. Suppose that there is such a solution $(x,y)$ satisfying $x\in(S_{11}\cup\{-1\})$. If $y\in S_{00}$, then (17) reduces to $y^{r}=(y+1)^{r}=-\frac{b}{2}$. This implies that $y^{\gcd(r,q-1)}=(y+1)^{\gcd(r,q-1)}$. If $\gcd(r,q-1)=1$, then we have $y=y+1$, a contradiction. If $\gcd(r,q-1)=2$, then we obtain $y^{2}=(y+1)^{2}$, which implies $y=-\frac{1}{2}$. Then, we get $\chi(y)=\chi\left(-\frac{1}{2}\right)=\chi\left(\frac{1}{2}\right)=\chi(y+1)=1$, a contradiction.
@@
Let $b\in\mathbb{F}_{q}^{*}$ and assume that $\gcd(r,q-1)\in\{1,2\}$. If $D_{00}(c)\leq 1$ for all $c\in\mathbb{F}_{q}^{*}$, then $B_{0000}(b)=0$.
@@
If $(x,y)\in S_{00}\times S_{00}$, then (17) leads to
$\begin{cases}x^{r}-y^{r}=\frac{b}{2},\\ (x+1)^{r}-(y+1)^{r}=\frac{b}{2}.\end{cases}$
Then we have $x^{r}-y^{r}=\frac{b}{2}=(x+1)^{r}-(y+1)^{r}$, which implies $(x+1)^{r}-x^{r}=(y+1)^{r}-y^{r}=c$ for some $c\in\mathbb{F}_{q}$. By Lemma 3.3, $D_{00}(0)=0$ and hence $c\neq 0$. Since $D_{00}(c)\leq 1$ for all $c\in\mathbb{F}_{q}^{*}$, we have $x=y$, and hence $b=0$, a contradiction. ∎
@@
The second equation of (18) has $\gcd(r,q-1)$ solutions. If $\gcd(r,q-1)=1$, then there exists a unique $x\in\mathbb{F}_{q}$ satisfying it. If $\gcd(r,q-1)=2$, then $r$ is even and hence $x^{r}=(-x)^{r}$ for all $x\in\mathbb{F}_{q}$. Thus, if $x=\theta$ is a solution of $(x+1)^{r}=\frac{b}{2}$, the other solution is $x=-\theta-2$. If both $\theta$ and $\theta-2$ lie in $S_{00}$, then $1=\chi(\theta+1)=\chi(-\theta-1)=-\chi(\theta+1)$, a contradiction. Therefore, the second equation of (18) admits at most one solution in $S_{00}$ in either case. For such $x$, the first equation of (18) has at most one solution $y$ with $\chi(y)=1$ in the same manner. Hence, $B_{0001}(b)\leq 1$, which completes the proof. ∎
@@
If (19) has a solution $y\in S_{01}$, then we have $2y^{r}=2=b$, which implies $y^{r}=-1$ and hence $r$ is odd. Then we obtain $\gcd(r,q-1)=1$ and y=1y=-1, which contradicts that $y\in S_{01}$.
@@
Combining Lemma 4.5 and Lemma 4.6 with the above equation, we have $\beta_{F_{r}}(1,b)\leq 2$, which completes the proof. ∎